An IT employee in Noida frequents one of the two restaurants KARIM or NAZEER, choosing Chicken's item 70% of the time and fish's item 30% of the time. Regardless of where he goes, he orders Afghani Chicken 60% of his visits.

a) The next time he goes into a a restaurants, what is the probability that he goes to KARIM and orders Afghani chicken.


b) Are the two events in part a independent? Explain.


c) If he goes into restaurants and orders Afghani chicken, what is the probability that he is at NAZEER.


d) what is the probability that he goes TO KARIM or orders Afghani chicken or both?


1
Expert's answer
2020-08-02T12:32:37-0400

A)P=0.7*0.6=0.42

B) Events are independent bevause he orders Afghani Chicken regardless of where he goes

C) P=30 %,  he orders Afghani Chicken regardless of where he goes

D) P=0.7*(1-0.6)+0.3*0.6+ 0.7*(0.6)=0.88